# Fuzzy Classification Aggregation

### v1

@@ Abstract
We consider the problem where a set of individuals has to classify $m$ objects into $p$ categories and does so by aggregating the individual classifications. We show that if $m\geq 3$, $m\geq p\geq 2$, and classifications are fuzzy, that is, objects belong to a category to a certain degree, then an independent aggregator rule that satisfies a weak unanimity condition belongs to the family of Weighted Arithmetic Means. We also obtain characterization results for $m=p=2$.
@@ 1 Introduction
A more general problem, where a group of individuals has to classify $m$ objects into $p$ different categories, has been considered by Maniquet and Mongin, (2016). In their work, they study the case where there are at least as many objects as categories, at least three categories and all the categories must be filled with at least one object. [^2] This surjectivity condition becomes relevant, for example, in situations where $m$ workers have to be assigned to $p$ tasks, and no task can be left unassigned. This aspect is combined with other three seemingly natural properties, namely Unanimity, which states that when all individuals make the same classification then the aggregate classification complies; Independence, which is the requirement that an object be classified by the aggregator in the same way at any two classification profiles if every individual classifies it equally in both profiles; and Non-Dictatorship, which requires that there is no individual such that her classification is always selected. The result is that there is no aggregator satisfying these three conditions. Recently, Cailloux et al., (2024) weaken the Unanimity axiom and find a weakening of the impossibility result that holds for $m\geq 3$ and $p\geq 2$, with the existence of an essential dictator. This is an individual such that a permutation of her classification is always selected. They also characterize the unanimous and independent aggregators when $m=p=2$.
@@
Our work considers the problem where $m$ objects have to be classified into $p$ categories, $m\geq p\geq 2$, and the classifications, both by the individuals and the rule, indicate degrees of membership of the objects to each of the categories. Consider a scenario where a national government allocates equal funds to all regions, with the stipulation that the central government determines how the funds are distributed. For instance, funds might be allocated for security, education, and health, ensuring that each category, across the country, receives at least the amount provided by the government to each region. As an illustration, one region could be tasked with allocating half of the funds to security, a quarter to education, and a quarter to health. Another scenario involves assigning individuals to various mandatory tasks with equal time requirements, where proportions dictate how each person’s time is allocated. For instance, if tasks include laundry, lawn mowing, and grocery shopping, Bob might spend half of his time on laundry, a quarter on mowing the lawn, and the remaining quarter on grocery shopping.
@@
We avoid both impossibility results, the crisp and the fuzzy ones, by considering a different, but also very natural, fuzzy surjectivity condition. Building on some previous results in functional analysis by Aczél and Wagner, (1980, 1981), and Wagner, (1982), we show that the only aggregators that are independent and that satisfy a weak version of Unanimity, are the Weighted Arithmetic Means.
Section 2 presents the basic notions and axioms that we use, while we present the results in Section 3. Finally, Section 4 contains some concluding remarks.
@@ 2 Basic Notions and Axioms
Let $N=\{1,\ldots,n\}$ be a finite set of agents and let $X=\{x_{1},\ldots,x_{m}\}$ be $m$ objects that need to be classified into the $p$ categories of a set $P$, with $m\geq p\geq 2$.
The individuals classify each object according to a partial degree of membership to each category. In the crisp setting, classifications are surjective mappings $c:X\rightarrow P$. In the fuzzy setting, a fuzzy classification is a mapping $c:X\rightarrow[0,1]^{p}$ such that $\sum_{j=1}^{m}c(x_{j})_{t}\geq 1$ for each $t\in\{1,\ldots,p\}$ and $\sum_{t=1}^{p}c(x)_{t}=1$ for all $x\in X$. The former condition is the fuzzy counterpart of the surjectivity of the classification function, while the latter is the fuzzy counterpart of the assumption that every object $x_{j}$ must be assigned to exactly one category by each voter. It is easy to see that when $m=p$, then all the inequalities turn into equalities.
We use $\mathcal{C}$ to denote the set of fuzzy classifications, and every $\bm{c}=(c_{1},\ldots,c_{n})\in\mathcal{C}^{N}$ is a fuzzy classification profile. Given $\mathbf{c}\in\mathcal{C}^{N}$ and $x\in X$, we denote with $\mathbf{c}^{x}\in[0,1]^{N\times P}$ the fuzzy classification profile restricted to x, such that the entry $\mathbf{c}^{x}_{ij}$ indicates the degree of membership of the object $x$ to the category $j$, according to the agent $i$. Thus, $\bm{c}^{x}_{ij}=c_{i}(x)_{j}$.
@@
$\alpha_{\bm{w}}(\bm{c})(x)=w_{1}c_{1}(x)+\ldots+w_{n}c_{n}(x).$
We say that the set of weights is degenerate if there is an $i\in N$ such that $w_{i}=1$. If we think of the FCAF as a group of experts classifying objects, this FCAF can be appropriate for situations where the individuals differ in expertise, with more experienced voters having higher weights. In particular, if $w_{i}=\frac{1}{n}$ for all $i\in N$, we call this FCAF the Arithmetic Mean.
@@
If we consider the Weighted Arithmetic Mean with a set of non-degenerate weights $\bm{w}=(\frac{1}{2},0,\frac{1}{2})$, we obtain the following social classifications:
@@
In the following, we introduce some axioms that an FCAF may satisfy. The first axiom states that the fuzzy social classification of an object in two different fuzzy classification profiles does not change if the classification regarding that object is the same in both profiles for every individual.
@@
Let $\bm{c,c^{\prime}}\in\mathcal{C}^{N}$ and $x\in X$ be such that $\bm{c}_{i}(x)=\bm{c^{\prime}}_{i}(x)$ for all $i\in N$. If an FCAF is independent, then $\alpha(\bm{c})(x)=\alpha(\bm{c^{\prime}})(x)$.
@@
An FCAF is symmetric if for all $x,y\in X$ and all $\bm{c}\in\mathcal{C}$ such that $\bm{c}_{i}(x)=\bm{c}_{i}(y)$ for all $i\in N$, it is the case that $\alpha(\bm{c})(x)=\alpha(\bm{c})(y)$.
Wagner, (1982) shows that a symmetric FCAF is also independent, and thus has the same elementary FCAF for every object. The next property states that if there is an object that is unanimously classified by the individuals, then the FCAF has to classify that object accordingly.
@@
An FCAF is zero unanimous if for all $\bm{c}\in\mathcal{C}$ where there is an $x\in X$ and a category $t\in P$ such that $c_{1}(x)_{t}=\cdots=c_{n}(x)_{t}=0$, it is the case that $\alpha(\bm{c})(x)_{t}=0$.
@@
An FCAF satisfies fuzzy consensus if $\alpha(\mathbf{c})(x)_{t}\in[\min_{i\in N}c_{i}(x)_{t},\max_{i\in N}c_{i}(x)_{t}]$, for every object $x\in X$ and for every category $t\in\{1,\ldots,p\}$.
@@
An FCAF is Non-Dictatorial if there is no individual $i\in N$ such that for all $x\in X$, it is the case that $\alpha(\mathbf{c})(x)=c_{i}(x)$.
@@
For a classification profile $\bm{c}\in\mathcal{C}$ and a permutation $\sigma:N\rightarrow N$ we define $\sigma(\bm{c})$ as the classification profile such that $\sigma(\bm{c})=\{\bm{c}_{\sigma(1)},\ldots,\bm{c}_{\sigma(n)}\}$. An FCAF is anonymous if $\alpha(\sigma(\bm{c}))=\alpha(\bm{c})$.
@@ 3 Results
Let $m\geq 3$ and $m\geq p\geq 2$. A Fuzzy Classification Aggregation Function satisfies Independence, Zero Unanimity, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
Now we present a result by Aczél and Wagner, (1981) that is useful for our proof.
@@
(Aczél and Wagner,, 1981) A family of mappings $\{\bm{c}^{j}:[0,1]^{N}\rightarrow[0,1]\}_{j=1}^{m}$ satisfies the following conditions:
  1. 1.
    $\bm{c}^{j}(0,\ldots,0)=0$ for $j=1,\ldots,m$, and
  2. 2.
    if $\sum_{j=1}^{m}(x_{j1},\ldots,x_{jn})=(1,\ldots,1)$, then $\sum_{j=1}^{m}\bm{c}^{j}(x_{j1},\ldots,x_{jn})=1$
if, and only if, there exists a set of weights $(w_{1},\ldots,w_{n})$ such that for $j=1,\ldots,m$ we have that
$\bm{c}^{j}(x_{j1},\ldots,x_{jn})=w_{1}x_{j1}+\cdots+w_{n}x_{jn}.$
@@
That a Weighted Arithmetic Mean with a non-degenerate set of weights satisfies the three axioms is straightforward to see. For the if part, we use the result by Aczél and Wagner, (1981). Let $\alpha$ be an FCAF that satisfies Independence and Zero Unanimity. Recall that an independent FCAF can be seen as a collection of elementary FCAFs. So we can consider a family of mappings $\{\alpha_{x_{j}}:[0,1]^{N\times P}\rightarrow[0,1]^{P}\}_{j=1}^{m}$. For a fixed category $t\in P$, we have that conditions $(1)$ and $(2)$ are satisfied by $(\alpha_{x_{j}})_{t}$.555$(\alpha_{x_{j}})_{t}$ is the projection of $\alpha_{x_{j}}$ over the coordinate $t$. Thus we have that for each category $t$ there is a set of weights $(w^{t}_{1},\ldots,w^{t}_{n})$ such that for $j=1,\ldots,m$, it is the case that666$(x_{ji})_{t}$ is the projection of $x_{ji}$ over the coordinate $t$.
$(\alpha_{x_{j}})_{t}((x_{j1})_{t},\ldots,(x_{jn})_{t})=w^{t}_{1}(x_{j1})_{t}+\cdots+w^{t}_{n}(x_{jn})_{t}.$
@@
Let $m\geq 3$ and $m\geq p\geq 2$. A Fuzzy Classification Aggregation Function satisfies Independence, Zero Unanimity, and Anonymity if, and only if, it is an Arithmetic Mean.
@@
Let $m\geq 3$ and $m\geq p\geq 2$. A Fuzzy Classification Aggregation Function satisfies Independence, Unanimity, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
@@
Let $m\geq 3$ and $m\geq p\geq 2$. A Fuzzy Classification Aggregation Function satisfies Independence, Fuzzy Consensus, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
@@
Let $m\geq 3$ and $m\geq p\geq 2$. A Fuzzy Classification Aggregation Function that satisfies Fuzzy Consensus, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
It is worth mentioning that the previous results are only valid for $m>2$. For the case of $2$ objects and $2$ categories, Independence is trivially satisfied (due to the surjectivity conditions). This forces us to use stronger axioms to obtain a characterization.
For symmetric and Zero Unanimous FCAFs, we can have FCAFs that are not Weighted Arithmetic Means, as the following result shows.
Theorem 3.
Let $m=p=2$. A symmetric Fuzzy Classification Aggregation Function $\alpha$ satisfies Zero Unanimity if, and only if, there is a function $h:[-\frac{1}{2},\frac{1}{2}]^{n}\rightarrow[-\frac{1}{2},\frac{1}{2}]$ where
$h(x_{1},\ldots,x_{n})=-h(-x_{1},\ldots,-x_{n}),$ (1)
@@
$h\left(\dfrac{1}{2},\ldots,\dfrac{1}{2}\right)=\dfrac{1}{2},$ (2)
@@
Before the proof, we present the folowing result by Wagner, (1982).
@@
(Wagner,, 1982) Let $m=p=2$. A symmetric mapping $\{\bm{c}:[0,1]^{N\times 2}\rightarrow[0,1]^{2}\}_{j=1}^{m}$ satisfies the following conditions:
  1. 1.
    $\bm{c}^{x_{j}}(0,\ldots,0)=0$ for $j=1,\ldots,m$, and
  2. 2.
    if $\sum_{j=1}^{2}(x_{j1},\ldots,x_{jn})=(1,\ldots,1)$, then $\sum_{j=1}^{2}\bm{c}^{x_{j}}(x_{j1},\ldots,x_{jn})=1$
if, and only if, there exists a function $h:[-\frac{1}{2},\frac{1}{2}]^{n}\rightarrow[-\frac{1}{2},\frac{1}{2}]$ where
$h(x_{1},\ldots,x_{n})=-h(-x_{1},\ldots,-x_{n}),$ (3)
@@
$h\left(\dfrac{1}{2},\ldots,\dfrac{1}{2}\right)=\dfrac{1}{2},$ (4)
such that for each $(x_{1},x_{2})\in[0,1]^{2}$,
$\bm{c}(x_{1},x_{2})=\left(h\left(x_{11}-\frac{1}{2},\ldots,x_{n1}-\frac{1}{2}\right)+\frac{1}{2},h\left(x_{12}-\frac{1}{2},\ldots,x_{n2}-\frac{1}{2}\right)+\frac{1}{2}\right).$
Proof of Theorem 3..
That such an FCAF satisfies Zero Unanimity is straightforward to see. For the if part, we use the result by Wagner, (1982). Thus, we have for each category, a mapping $h$ satisfying Wagner Theorem’s conditions, namely $h^{1}$ and $h^{2}$. What is left to see is that $h^{1}=h^{2}$. We have the following conditions that must be satisfied by $\alpha(\bm{c})(x_{1})$ and $\alpha(\bm{c})(x_{2})$:

### v3

@@ Abstract
We consider the problem where a set of individuals has to classify $m$ objects into $p$ categories and does so by aggregating the individual classifications. We show that if $m\geq 3$, $m\geq p\geq 2$, and classifications are fuzzy, that is, objects belong to a category to a certain degree, then an optimal and independent aggregator rule that satisfies a weak unanimity condition belongs to the family of Weighted Arithmetic Means. We also obtain characterization results for $m=p=2$.
@@ 1 Introduction
A more general problem, where a group of individuals has to classify $m$ objects into $p$ different categories, has been considered by Maniquet and Mongin, (2016). In their work, they study the case where there are at least as many objects as categories, at least three categories and all the categories must be filled with at least one object. [^2] This surjectivity condition becomes relevant, for example, in situations where $m$ workers have to be assigned to $p$ tasks, and no task can be left unassigned. It can also be seen as a sanity check of the classification process. For instance, consider a set of neural networks that is being trained to classify images into images of cats, dogs and rabbits. If a database of a million pictures with images of the three animals is used, then a neural network that fails to classify at least one image as a rabbit can be considered that is not reliable. This aspect is combined with three other seemingly natural properties, namely Unanimity, which states that when all individuals make the same classification then the aggregate classification complies; Independence, which is the requirement that an object be classified by the aggregator in the same way at any two classification profiles if every individual classifies it equally in both profiles; and Non-Dictatorship, which requires that there is no individual such that her classification is always selected. The result is that there is no aggregator satisfying these three conditions. Recently, Cailloux et al., (2024) weaken the Unanimity axiom and find a weakening of the impossibility result that holds for $m>p\geq 2$, with the existence of an essential dictator. This is an individual such that a permutation of her classification is always selected. They also characterize the unanimous and independent aggregators when $m=p=2$.
The surjectivity condition in Maniquet and Mongin, (2016), turns out to be a rather demanding requirement. Without it, the impossibility result no longer holds, and many non-dictatorial aggregators satisfying Unanimity and Independence can be found in the literature (See, for example, Kasher and Rubinstein,, 1997).
@@
Our work considers the problem where $m$ objects have to be classified into $p$ categories, $m\geq p\geq 2$, and the classifications, both by the individuals and the rule, indicate degrees of membership of the objects to each of the categories. Two possible scenarios where this setting may arise, are the following. A national government has to allocate equal funds to all regions, to be used for different regional ministries (security, education, and health), with the central government providing some conditions on how the funds are distributed. It is natural to assume in this context, that the government sets a minimum amount of money that should be spent in each of the ministries, at least at a national level. Members of the government should decide on how the funds might be allocated for security, education, and health, ensuring that each category, across the country, receives at least the amount provided by the government to each region. As an illustration, one region could be tasked with allocating half of the funds to security, a quarter to education, and a quarter to health. Another scenario involves assigning individuals to various mandatory tasks with equal time requirements, where proportions dictate how each person’s time is allocated. For instance, if tasks include laundry, lawn mowing, and grocery shopping, Bob might spend half of his time on laundry, a quarter on mowing the lawn, and the remaining quarter on grocery shopping. [^3] These examples illustrate why the notion of “degree of membership” can be considered natural to use, instead of considering probabilities. For instance, in the tasks allocation example, it is better if we know that Bob will spend with certainty half of his time doing the laundry, rather than knowing that there is a $50\%$ chance of Bob leaving all the clothes dirty.
@@
We avoid both impossibility results, the crisp and the fuzzy ones, by considering a different, but also very natural, fuzzy surjectivity condition. This condition, which in the crisp setting is extremely demanding, becomes a requirement that in the fuzzy setting allows the existence of a family of aggregators satisfying certain particular properties. Thus, it is demanding enough to find a characterization result, but not so much to not be able to find one. Building on some previous results in functional analysis by Aczél and Wagner, (1980); Aczél et al., (1984), and Wagner, (1982), we show that the only optimal aggregators that are independent and that satisfy a weak version of Unanimity, are the Weighted Arithmetic Means.
Considering different settings for already established problems is a common technique used in social choice to avoid impossibility results. The result by Gibbard, (1973) and Satterthwaite, (1975), that states in the deterministic setting that the only rule immune to manipulations is the dictatorial one, is avoided when the outcomes of the rule are lotteries over the set of alternatives (Gibbard,, 1977). We consider that our paper is related to Maniquet and Mongin, (2016)’s paper in an analogous way as Dutta, (1987) and Billot, (1991) results are related to the Arrow,’s (1951) impossibility theorem. In their work, they consider a fuzzy setting and adapt Arrow’s theorems to it to obtain possibility results. Dutta, (1987) characterizes the oligarchic aggregators, while Billot, (1991) does the same for the leximin aggregator. The use of a fuzzy setting, that allows many interpretations of axioms that are used in crisp settings, proves advantageous for circumventing impossibility results.
The plan of the paper is as follows. Section 2 presents the basic notions and axioms that we use, while we present the results in Section 3. Finally, Section 4 contains some concluding remarks.
@@ 2 Basic Notions and Axioms
Let $N=\{1,\ldots,n\}$ be a finite set of individuals and let $X=\{x_{1},\ldots,x_{m}\}$ be $m$ objects that need to be classified into the $p$ categories of a set $P$, with $p\geq 2$.
The individuals classify each object according to a partial degree of membership to each category. In the crisp setting, introduced by Maniquet and Mongin, (2016), classifications are surjective mappings $c:X\rightarrow P$, that is, every category must have at least one object classified into. In the fuzzy setting, a fuzzy classification is a mapping $c:X\rightarrow[0,1]^{p}$ such that $\sum_{j=1}^{m}c(x_{j})_{t}\geq 1$ for each $t\in\{1,\ldots,p\}$ and $\sum_{t=1}^{p}c(x)_{t}=1$ for all $x\in X$. The former condition is the fuzzy counterpart of the surjectivity of the classification function, while the latter is the fuzzy counterpart of the assumption that every object $x$ must be assigned to exactly one category by each voter. It is easy to see that these conditions imply that there must be at least as many objects as categories, thus $m\geq p$.
We use $\mathcal{C}$ to denote the set of fuzzy classifications, and every $\bm{c}=(c_{1},\ldots,c_{n})\in\mathcal{C}^{N}$ is a fuzzy classification profile. Given $\mathbf{c}\in\mathcal{C}^{N}$ and $x\in X$, we denote with $\mathbf{c}^{x}\in[0,1]^{N\times P}$ the fuzzy classification profile restricted to x, such that the entry $\mathbf{c}^{x}_{ij}$ indicates the degree of membership of the object $x$ to the category $j$, according to the agent $i$. Thus, we have that $\bm{c}^{x}_{ij}=c_{i}(x)_{j}$. This allows the representation of a fuzzy classification profile as a set of $m$ matrices in $[0,1]^{N\times P}$, one for each object, where row $i$ indicates the classifications given by the individual $i$, and column $t$ of the matrix indicates the classifications given by the set of individuals into the category $t$.
@@
$\alpha_{\bm{w}}(\bm{c})(x)=w_{1}c_{1}(x)+\cdots+w_{n}c_{n}(x)$
We say that the set of weights is degenerate if there is an $i\in N$ such that $w_{i}=1$. If we think of the FCAF as a group of experts classifying objects, this FCAF can be appropriate for situations where the individuals differ in expertise, with more experienced voters having higher weights. [^4] Weighted aggregators have been previously studied in the social choice literature, in particular in the setting where monetary transfers are allowed (see Roberts,, 1979; Mishra and Sen,, 2012). See also Stone, (1961) and Dietrich and List, (2016) for an use of this aggregator in opinion pooling. In particular, if $w_{i}=\frac{1}{n}$ for all $i\in N$, we call this FCAF the Arithmetic Mean.
@@
In this example, individual $2$ classifies the object $x_{2}$ as being only part of category $3$ (second row of the second matrix). If we consider the Weighted Arithmetic Mean with a non-degenerate set of weights $\bm{w}=(\frac{1}{2},0,\frac{1}{2})$, we obtain the following social classifications:
@@
In the following, we introduce a number of axioms, i.e., fundamental normative requirements that, depending on the specific classification problem at hand, a reasonable FCAF should satisfy. The first axiom can be seen as an efficiency property. It states that if all the individuals agree on how many objects should be classified into a given category, then the social classification must respect that.
Definition 1 (Optimality).
An FCAF is optimal if for all $\bm{c}\in\mathcal{C}^{N}$ such that $\sum_{j=1}^{m}c_{i}(x_{j})_{t}=h$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha(\bm{c})(x_{j})_{t}=h$.
This axiom is useful in a setting where categories need to fulfil certain requirements and resources should not be wasted. For example, in the case where a set of managers has to classify a group of workers as members of different departments within a company, in order to make them operative. If all the managers consider that department A requires only two workers to be operative, then an optimal aggregator only classifies two workers into that department. The following axiom states that the fuzzy social classification of an object in two different fuzzy classification profiles does not change if the classification regarding that object is the same in both profiles for every individual.
@@
An FCAF is independent if given $\bm{c,c^{\prime}}\in\mathcal{C}^{N}$ and $x\in X$ such that $\bm{c}_{i}(x)=\bm{c^{\prime}}_{i}(x)$ for all $i\in N$, it is the case that $\alpha(\bm{c})(x)=\alpha(\bm{c^{\prime}})(x)$.
@@
An FCAF is symmetric if for all $x,y\in X$ and all $\bm{c}\in\mathcal{C}^{N}$ such that $\bm{c}_{i}(x)=\bm{c}_{i}(y)$ for all $i\in N$, it is the case that $\alpha(\bm{c})(x)=\alpha(\bm{c})(y)$.
Wagner, (1982) shows that an optimal and symmetric FCAF is also independent, and thus has the same elementary FCAF for every object. The next property states that if there is an object that is unanimously classified by the individuals, then the FCAF has to classify that object accordingly.
@@
An FCAF is zero unanimous if for all $\bm{c}\in\mathcal{C}^{N}$ where there is an $x\in X$ and a category $t\in P$ such that $c_{1}(x)_{t}=\cdots=c_{n}(x)_{t}=0$, it is the case that $\alpha(\bm{c})(x)_{t}=0$.
@@
An FCAF satisfies Fuzzy Consensus if for every object $x\in X$ and for every category $t\in\{1,\ldots,p\}$, it is the case that $\alpha(\mathbf{c})(x)_{t}\in[\min_{i\in N}c_{i}(x)_{t},\max_{i\in N}c_{i}(x)_{t}]$.
@@
An FCAF is non-dictatorial if there is no individual $i\in N$ such that for all $x\in X$ and all $\bm{c}\in\mathcal{C}^{N}$, it is the case that $\alpha(\mathbf{c})(x)=c_{i}(x)$.
@@
For a classification profile $\bm{c}\in\mathcal{C}^{N}$ and a permutation $\sigma:N\rightarrow N$ we define $\sigma(\bm{c})$ as the classification profile such that $\sigma(\bm{c})=\{\bm{c}_{\sigma(1)},\ldots,\bm{c}_{\sigma(n)}\}$. An FCAF is anonymous if for all $\bm{c}\in\mathcal{C}^{N}$ and all permutations $\sigma:N\rightarrow N$, it is the case that $\alpha(\sigma(\bm{c}))=\alpha(\bm{c})$.
@@ 3 Results
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Zero Unanimity, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
Now we present a result by Aczél et al., (1984) that is useful for our proof.
@@
(Aczél et al.,, 1984) A family of mappings $\{\alpha^{j}:[0,1]^{N}\rightarrow[0,1]\}_{j=1}^{m}$ satisfies Zero Unanimity and the following condition:
$\text{if }z^{j}\in[0,1]^{N}\text{ and }\sum_{j=1}^{m}z^{j}=(1,\ldots,1),\text{then }\sum_{j=1}^{m}\alpha^{j}(z^{j})=1$ (1)
if, and only if, there exists a set of weights $(w_{1},\ldots,w_{n})$ such that for $j=1,\ldots,m$ and any $z\in[0,1]^{N}$, we have that
$\alpha^{j}(z)=w_{1}z_{1}+\cdots+w_{n}z_{n}.$
@@
That a Weighted Arithmetic Mean with a non-degenerate set of weights satisfies the four axioms is straightforward to see. For the only if part, we use the result by Aczél et al., (1984). Let $\alpha$ be an optimal FCAF that satisfies Independence and Zero Unanimity. Recall that an independent FCAF can be seen as a collection of elementary FCAFs. So we can consider a family of mappings $\{\alpha_{x_{j}}:[0,1]^{N\times P}\rightarrow[0,1]^{P}\}_{j=1}^{m}$. By optimality, we obtain that if $\sum_{j=1}^{m}c_{i}(x_{j})_{t}=h$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha_{x_{j}}(\bm{c}^{x_{j}})_{t}=h$. Then we have that the family of mappings $\{(\alpha_{x_{j}})_{t}:[0,1]^{N}\rightarrow[0,1]\}_{j=1}^{m}$ satisfies condition (1) for every $t\in P$777$(\alpha_{x_{j}})_{t}$ is the projection of $\alpha_{x_{j}}$ over the category $t$. Thus we have that for each category $t$ there is a set of weights $(w^{t}_{1},\ldots,w^{t}_{n})$ such that for $j=1,\ldots,m$, it is the case that
$(\alpha_{x_{j}})_{t}(\bm{c}^{x_{j}}_{1t},\ldots,\bm{c}^{x_{j}}_{nt})=w^{t}_{1}\bm{c}^{x_{j}}_{1t}+\cdots+w^{t}_{n}\bm{c}^{x_{j}}_{1t}.$
@@
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Zero Unanimity, and Anonymity if, and only if, it is an Arithmetic Mean.
@@
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Unanimity, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
@@
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Fuzzy Consensus, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
@@
[…]
It is worth mentioning that the previous results are only valid for $m\geq 2$. For the case of $m=2$ objects and $p=2$ categories, Independence is trivially satisfied (due to the surjectivity conditions). This forces us to use a stronger axiom to obtain a characterization.
For optimal, symmetric and zero unanimous FCAFs, we can have FCAFs that are not Weighted Arithmetic Means, as the following result shows.
Proposition 1.
Let $m=p=2$. An optimal and symmetric Fuzzy Classification Aggregation Function $\alpha$ satisfies Zero Unanimity if, and only if, there is a function $h:[-\frac{1}{2},\frac{1}{2}]^{n}\rightarrow[-\frac{1}{2},\frac{1}{2}]$ where
$h(x_{1},\ldots,x_{n})=-h(-x_{1},\ldots,-x_{n}),$
@@
$h\left(\dfrac{1}{2},\ldots,\dfrac{1}{2}\right)=\dfrac{1}{2},$
@@
Before the proof, we present the following result by Wagner, (1982).
@@
(Wagner,, 1982) Let $m=p=2$. An optimal and symmetric mapping $\alpha:[0,1]^{N\times 2}\rightarrow[0,1]^{2}$ satisfies Zero Unanimity if, and only if, there exists a function $h:[-\frac{1}{2},\frac{1}{2}]^{n}\rightarrow[-\frac{1}{2},\frac{1}{2}]$ where
$h(y_{1},\ldots,y_{n})=-h(-y_{1},\ldots,-y_{n}),$
@@
$h\left(\dfrac{1}{2},\ldots,\dfrac{1}{2}\right)=\dfrac{1}{2},$
such that for each $\bm{c}\in[0,1]^{N\times 2}$, we have that
$\alpha(\bm{c})=\left(h\left(\bm{c}_{11}-\frac{1}{2},\ldots,\bm{c}_{n1}-\frac{1}{2}\right)+\frac{1}{2},h\left(\bm{c}_{12}-\frac{1}{2},\ldots,\bm{c}_{n2}-\frac{1}{2}\right)+\frac{1}{2}\right).$
Proof of Proposition 1..
That such an FCAF satisfies Zero Unanimity is straightforward to see. For the if part, we use the result by Wagner, (1982). By Optimality and Zero Unanimity, for each category the hypotheses of Wagner, (1982) are satisfied and thus, we have a mapping $h$ satisfying Wagner,’s conditions, namely $h^{1}$ and $h^{2}$. What is left to see is that $h^{1}=h^{2}$. We have the following conditions that must be satisfied by $\alpha(\bm{c})(x_{1})$ and $\alpha(\bm{c})(x_{2})$: